# Synthesizing data products, mathematical models, and observational measurements for lake temperature forecasting

### v1

@@ 3.2 Heteroscedastic surrogate modeling
The first way that we make the surrogate modeling enterprise more manageable is to disregard the ensemble column $e$ in $\mathbf{X}_{\_}{n_{\_}M}$ and thereby treat forecasts from each $e$ as independent conditional on time and depth. This statistical modeling assertion is purely for computational convenience; it is not faithful to the simulation mechanism. While NOAA ensemble members are independent of one another, each realization evolves in a time-dependent manner. Forecasts from the prior day or depth for a specific ensemble member might contain valuable data for predicting the following day or depth for that same member, and potentially others. Nonetheless, our assumption of independence aligns with the essence of ensemble modeling. The 31-member NOAA ensemble allows for the measurement of a variety of environmental conditions, whose variability grows over time. We apply these to GLM simulations, as depicted in Figure 3.
This “modeling hack” means that each of 31 ensemble forecasts may be treated as replicates. Baker et al. (2022) explain that replicates bring two kinds of efficiency. One is computational: the number of sufficient statistics are fewer – 31-fold fewer in this instance. The other is statistical: replicates allow a pure glimpse at variance, without needing to first separate signal from noise. This is especially important when noise levels are changing along with the mean, in a potentially nonlinear way, as we have in Figure 3. The most effective methods for modeling such nonlinear heteroskedasticity (e.g., Goldberg, Williams and Bishop, 1997; Ankenman, Nelson and Staum, 2008; Quadrianto et al., 2009; Lázaro-Gredilla and Titsias, 2011; Binois, Gramacy and Ludkovski, 2018) leverage replication. Stochastic kriging (SK; Ankenman, Nelson and Staum, 2008), one of the first of these, is ideal for our setting. SK is sometimes criticized for (a) being moment-based (rather than being completely likelihood-based); and (b) for not jointly modeling the mean and variance process(es). These are valid concerns when the degree of replication is low. Ankenman, Nelson and Staum (2008) recommend at least ten replicates for each otherwise unique input. We have 31. Joint, likelihood-based modeling, as suggested by many of the other references cited above, is better statistically but much more work computationally.
Suppose that simulation outputs $\mathbf{Y}_{\_}{n_{\_}M}$ correspond to $n\ll n_{\_}M$ unique inputs in $\mathbf{X}_{\_}{n_{\_}M}$ when ignoring the $e$ column. In our case $n=n_{\_}M/31$. Let $\bar{\mathbf{X}}_{\_}n$ denote the matrix of those unique inputs, again without the $e$ column, so that $\bar{\mathbf{X}}_{\_}n$ is $n\times 4$. Now, consider each row $\bar{\mathbf{x}}_{\_}i$ of $\bar{\mathbf{X}}_{\_}n$, and let $y_{\_}j(\bar{\mathbf{x}}_{\_}i)$ denote the $j^{\mathrm{th}}$ replicate associated with $\bar{\mathbf{x}}_{\_}i$ among the $n_{\_}i$ replicates in $\mathbf{X}_{\_}{n_{\_}M}$. We have that all $n_{\_}i=31$. Then, calculate the first two moments for these replicates as
@@
Finally, store these in $\bar{\mathbf{Y}}_{\_}n$ and $\bar{\mathbf{S}}_{\_}n$, respectively, to line up with the rows of $\bar{\mathbf{X}}_{\_}n$. Note that we use a bar on top of $\bar{\mathbf{X}}_{\_}n$ not to indicate that any averaging is involved in their calculation, but instead to visually match up with $\bar{\mathbf{Y}}_{\_}n$ and $\bar{\mathbf{S}}_{\_}n$, which do involve averaging. Also note that we are dropping the $M$ subscripts to streamline the notation here with the understanding that we are only using this on the computer model runs. If the field data also had replicates we could additionally make use of them there, but our sensors record only one measurement per day.
SK involves modeling $\bar{\mathbf{Y}}_{\_}n$ and $\bar{\mathbf{S}}_{\_}n$ via GPs rather than operating directly on the original $D_{\_}{n_{\_}M}$. It turns out that these $2n$ quantities are not sufficient for $\mathbf{Y}_{\_}{n_{\_}M}$ in the context of prediction, but they are close. Binois, Gramacy and Ludkovski (2018) show that there are $12n+$ 1 sufficient statistics. Nevertheless, Ankenman, Nelson and Staum (2008) argue that inference based on these $2n$ quantities is unbiased asymptotically and minimizes mean-squared error.
Suppose we wish to predict $\bar{Y}$ at new inputs $\mathcal{X}$ with a GP. If the degree of replication is uniform, e.g., $n_{\_}i=31$ for all $i$, then we could simply plug the $\bar{\mathbf{Y}}_{\_}n$ values, along with $\bar{\mathbf{X}}_{\_}n$, into Eq. (4) after solving for MLE hyperparameters. We can do the same thing for the second moments $\bar{\mathbf{S}}_{\_}n$. This is the essence of SK with two caveats: (a) uniform replication is rare, although that is the case for NOAA-GLM; (b) it helps to explicitly link the mean and variance processes when making forecasts. First, fit a GP to $\bar{\mathbf{S}}_{\_}n$, often on a transformed scale (e.g., via log or a log/square-root hybrid (Johnson et al., 2018)) in order to respect positivity. Specifically, fit $\mathcal{GP}(D^{(v)}_{\_}n)$ where $D^{(v)}_{\_}n=(\bar{\mathbf{X}}_{\_}n,T(\bar{\mathbf{S}}_{\_}n))$ via MLE. For $T(\cdot)$ we prefer the square root to model standard deviations, although a logarithm also works well. Then feed the unique training $\bar{\mathbf{X}_{\_}n}$ values into the predictive equations (4), i.e., with $\mathcal{X}=\bar{\mathbf{X}}_{\_}n$. Let
@@
Similarly, obtain predictions at any new/forecasting coordinates $\mathcal{X}$ that may be of interest, like for a particular day $t$ and multiple horizons $h$ into the future. These $\mathcal{X}$s could represent genuine forecasts, or hindcasts that would be compared later to sensor measurements, which we discuss further in Section 3.4. Let $s(\mathcal{X})=\mathrm{Diag}(T^{-1}(\mu^{(v)}(\mathcal{X})))$ capture these back-transformed variance predictions – they are the predictive analog of Eq. (8). Using these, and similar quantities described above, the SK equations are
$\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ $=\tau^{2}k(\bar{\mathbf{X}}_{\_}n,\mathcal{X})^{\top}(\tau^{2}k(\bar{\mathbf{X}}_{\_}n)+\mathbf{S}_{\_}n)^{-1}\boldsymbol{\bar{Y}}_{\_}{n}^{M}$ (9)
@@
Observe in Eq. (9) that augmenting the diagonal of $k(\bar{\mathbf{X}}_{\_}n)$ with $\mathbf{S}_{\_}n$ results in a scaled nugget-like term, by analogy to $\tau^{2}g$ in Eq. (4). It facilitates a signal-to-noise trade-off that is unique to each training data element and/or its degree of replication. Having $\tau^{2}$ implicitly a part of $\mathbf{S}_{\_}n$ prevents a cancellation of $\tau^{2}$ that would otherwise render a predictive mean equation that is free of this quantity. Smoothed (squared) standard errors $\mathbf{S}_{\_}n$ naturally downweight large variance/low replication training data. Crucially, it may be shown that $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})=\mu_{\_}{n_{\_}M}(\mathcal{X})$ from Eq. (4), identically, i.e., the full $n_{\_}M$-data predictive mean can be recovered from GP smoothing over the replicate-wise $2n$ almost-sufficient statistics, which could represent a huge computational savings when $n\ll n_{\_}M$ as amplified by cubic matrix decomposition.
@@
Figure 5: 2d slices of predictive mean $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ (left) and variance $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}$ (right) from the (heteroscedastic) GLM surrogate over depth (m) and horizon (days). Together, these summarize a hindcast for 2022-02-19.
@@
To illustrate, consider Figure 5 which is derived from fits of two GPs, for mean and (square root) variance, to a simulation campaign spanning three years over a thirty-day horizon. The data represented in the plot comprises a 2d slice of predictions from those two fits corresponding to a forecast beginning on 2022-02-19, offering a surrogate for the NOAA-GLM hindcast beginning on that date. Focusing first on the left panel, showing $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ from Eq. (9), observe that temperature is decreasing with increasing depth, and also generally increasing over horizon, most prominently at the surface (low depths). Both relationships are clearly nonlinear. Now looking at the right panel providing $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}=\mathrm{Diag}(\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X}))$ we see a similar relationship in the spread of possible temperatures. Variance is much lower at lower depths, even at longer horizons, which makes sense because these depths are more insulated from weather at the surface. At the surface, the variance grows quickly farther out into the future.
Figure 6: GLM ensemble forecasts 1–30 day-ahead (gray lines) beginning on Feb 10, 2021 along with a 1d slice of the SK surrogate (solid orange line) 90% CIs and PI (dotted and dashed orange lines), respectively.
Figure 6 shows a 1d slice of the surface(s) in Figure 5 at 1m depth along with additional information. Gray lines indicate raw NOAA-GLM simulations (i.e., a subset of the training data $D_{\_}{n_{\_}M}$). Filled purple dots indicate field/sensor observations at 1m depth on the days in question. These latter values are not part of the analysis at this time and are provided for reference only. Of particular note are the dashed orange error-bars indicating a 90% PI derived from $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ and $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}$. Observe that the 90% PIs have appropriate coverage. Dotted lines indicate 90% CIs on the mean that are derived by eliminating the $s(\mathcal{X})$ term in Eq. (10), as an analogue to Eq. (5).
It is this squared standard error which is crucial to our bias-correcting analysis described in Section 3.4. Observe that, at least in this view, NOAA-GLM forecasts and their surrogate have a cold bias compared to the field data. But first we must describe how we make these SK calculations with large training data sizes. Although reducing $n_{\_}M$ from around 8 million to $n$ around 270,000 (900 reference dates $\times$ 10 depths $\times$ 30 horizons) is helpful, this $n$ is still too large for conventional GP fitting.
@@ 3.3 GP surrogates and scalability
The surrogates behind the views in Figures 5 and 6 were constructed using the approximate methods we describe here. Scalable GP approximation has been an active field of late. Techniques include compactly supported kernels (Gneiting, 2002; Kaufman et al., 2011), fixed-rank kriging (Cressie and Johannesson, 2008) inducing points (Banerjee et al., 2008; Quinonero-Candela and Rasmussen, 2005; Hensman, Fusi and Lawrence, 2013) local divide-and-conquer (Emery, 2009; Gramacy, 2016) and the Vecchia/nearest-neighbor approximation (Datta et al., 2016; Wu, Pleiss and Cunningham, 2022; Vecchia, 1988; Katzfuss and Guinness, 2021; Guinness, 2018; Zhang and Katzfuss, 2022; Katzfuss et al., 2020; Katzfuss, Guinness and Lawrence, 2022; Katzfuss and Guinness, 2021). We have identified that the so-called scaled Vecchia approximation (SVecchia Katzfuss, Guinness and Lawrence, 2022) is particularly well-suited to our setting, and to surrogate modeling at large because it offers a global model, accommodates input dimension larger than two, allows coordinate-wise lengthscale estimation, and scales in flops that are quasi-linear in training data size, $n$. Crucially, a user-friendly open-source implementation is provided for R: https://github.com/katzfuss-group/scaledVecchia.
The crux of the Vecchia approximation is a simple identity linking joint and conditional distributions. Consider a vector $\mathbf{Y}_{\_}n$ of responses, like one filled with $\bar{y}_{\_}i$-values in Section 3.2 above. The joint probability of $\mathbf{Y}_{\_}n$ can be factorized as
Some additional implementation details make things even faster and allow for distributed computation, which we leave to our references (Stein, Chi and Welty, 2004; Guinness, 2018; Katzfuss and Guinness, 2021; Stroud, Stein and Lysen, 2017; Datta et al., 2016; Wu, Pleiss and Cunningham, 2022; Katzfuss et al., 2020). We have been satisfied with the performance of the defaults offered by the R library linked above, synthesizing many of those elements including composition of the $c(i)$. We settled on $m=30$ by entertaining an out-of-sample prediction exercise similar to that reported by Sauer, Cooper and Gramacy (2023), which showed diminished returns for larger $m$-values. Perhaps the most important aspect of that software, compared to other methods based on the Vecchia approximation, is that the effect of an input-dependent lengthscale $\boldsymbol{\gamma}$ can be learned along with scale $\tau^{2}$ and (scalar) nugget $g$ via Fisher scoring (Guinness, 2021). The authors call this the “Scaled Vecchia” approximation; however we’ll simply refer to it as Vecchia going forward with the understanding that input-dependent lengthscales are being learned. With this setup we are able to fit GPs with $n\approx 270{,}000$ as described in Section 3.2 in about twenty seconds on an ordinary workstation (64-bit Intel(R) Core(TM) i9-9900K CPU @ 3.60GHz with 16 cores).
@@
The Vecchia approximation has never been used, and nor to our knowledge has any other large-scale GP approximation, to model input-dependent variances, like through SK (Section 3.2). Fitting second moments $\mathcal{GP}(D_{\_}n^{(v)}$) in order to get a predictor for $T(\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathcal{X}})\})$ is straightforward, because this amounts to feeding in (transformed) $s_{\_}i$-values (7) as responses along with unique inputs $\bar{\mathbf{x}}_{\_}i$. The visual in the right panel of Figure 5 shows $\mu^{(v)}_{\_}n(\mathcal{X})$. Using Vecchia to fit the mean GP conditional on those variances (left panel of Figure 5) is harder because the R library does not allow the user to specify a vector of nuggets, which comprises $\mathbf{S}_{\_}n$ in Eqs. (910). We inspected the source code to determine if modifications could accommodate $\mathbf{S}_{\_}n$ but ultimately ended up implementing the following workaround instead.
@@
We begin by fitting first moments with an otherwise ordinary Vecchia approximated GP, i.e., $\mathcal{GP}(D_{\_}n^{(m)})$, where $D_{\_}n^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$. Let $\mu^{(m)}_{\_}n(\mathcal{X})$ denote the predictive mean of $\mathcal{GP}(D_{\_}n^{(m)})$. Note that $\mu^{(m)}_{\_}n\neq\mu^{\mathrm{SK}}_{\_}n$ from Eq. (9) because the former utilizes a covariance structure with a single nugget, i.e., $\tau^{2}(k(\bar{\mathbf{X}}_{\_}n)+g\mathbb{I}_{\_}n)$ following Eq. (4), and the latter uses $\tau^{2}k(\bar{\mathbf{X}}_{\_}n)+\mathbf{S}_{\_}n$. However, they are equal in expectation because GPs furnish unbiased predictors (see, e.g., Santner et al., 2018). Nevertheless the two can be quite different if the degree of replication is high because of the $\mathbf{n}$-vector in the denominator of Eq. (8). Specifically, if the replication degree varies significantly from one location to another, for instance $n_{\_}i\gg n_{\_}j$ for $\bar{\mathbf{x}}_{\_}i$ and $\bar{\mathbf{x}}_{\_}j$, then the predictions from $\mu^{\mathrm{SK}}_{\_}n(\mathbf{x})$ for $\mathbf{x}$ in the vicinity of $\bar{\mathbf{x}}_{\_}j$, the site with fewer replicates, could be considerably more precise than $\mu^{(m)}_{\_}n(\mathbf{x})$, which does not know to place less trust in $\bar{y}_{\_}j$ than in other locations, such as near $\bar{\mathbf{x}}_{\_}i$. However, in our setting we have uniform replication, with all $n_{\_}i=31$. In a study described momentarily, and summarized in Figure 7, we find there is essentially no difference between $\mu^{(m)}_{\_}n$, via Vecchia, and $\mu^{\mathrm{SK}}_{\_}n$.
@@
Approximating what SK provides for the variance of the first moment, $\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})$ in Eq. (10), with Vecchia, follows a similar thread. Whereas our development of the mean leverages the unbiased nature of Gaussian expectations, second moments follow standard calculations of Gaussian variances. Crucially, we do not need the entire covariance structure $\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})$ for bias correction in Section 3.4; we only need the diagonal. Those “variances of the mean” are Gaussian standard errors, i.e., $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}/n$, again leveraging uniform replication: $n_{\_}i=31$. Our Vecchia fit to second moments provides an estimate of $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}$ so all we need to do is divide this quantity by $n_{\_}i$, for any $i$, to get an estimate of the variance of the mean. However, we prefer a more conservative application of this in an SK-approximating context to avoid under-covering out-of-sample. Let $\mu^{(v)95}_{\_}n(\mathbf{x})$ denote the upper $95^{\mathrm{th}}$ quantile of predictive equations derived from $\mathcal{GP}(D_{\_}n^{(v)})$. We then combine this estimate with our first-moments’ fit, to obtain the following Vecchia approximation to SK.
$\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathbf{x})$ $=\mu_{\_}n^{(m)}(\mathbf{x})$ (12)
$\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ $=\mu^{(v)95}_{\_}n(\mathbf{x})/n_{\_}i+\sigma^{2(m)}_{\_}n(\mathbf{x}).$

### v2

@@ 3.2 Heteroscedastic surrogate modeling
We begin by discarding the ensemble column $\xi$ in $\mathbf{X}_{\_}{n_{\_}M}$ and treating forecasts from each $\xi$ as independent conditional on time and depth. This modeling assertion is purely for computational convenience; it is not faithful to the simulation mechanism. While NOAA ensemble members are independent of one another, each realization evolves in a time-dependent manner. Forecasts from the prior day or depth for a specific ensemble member might contain valuable data for predicting the following day or depth for that same member, and potentially others. Nonetheless, independence aligns with the essence of ensemble modeling (Thomas et al., 2020).
This “modeling hack” means that each of 31 ensemble forecasts may be treated as replicates. Baker et al. (2022) explain that replicates bring two kinds of efficiency. One is computational: the number of sufficient statistics are fewer – 31-fold fewer in this instance. The other is statistical: replicates allow a pure glimpse at variance, without needing to first separate signal from noise. This is especially important when noise levels are changing along with the mean, in a potentially nonlinear way as in Figure 3. The most effective methods for modeling such nonlinear heteroskedasticity (e.g., Goldberg, Williams and Bishop, 1997; Ankenman, Nelson and Staum, 2008; Quadrianto et al., 2009; Lázaro-Gredilla and Titsias, 2011; Binois, Gramacy and Ludkovski, 2018) leverage replication. Stochastic kriging (SK; Ankenman, Nelson and Staum, 2008), one of the first of these, is ideal for our setting, in part because it cuts corners relative to other methods [cited above]; e.g., (a) being moment-based (rather than likelihood-based); and (b) not jointly modeling the mean and variance process(es). Indeed, both are valid concerns when the degree of replication is low. Ankenman, Nelson and Staum (2008) recommend at least ten replicates for each otherwise unique input. We have 31, more than plenty on that account. Perhaps more importantly, SK is uniquely situated to interface with our second layer of approximation in Section 3.3.
Suppose that simulation outputs $\mathbf{Y}_{\_}{n_{\_}M}$ correspond to $n\ll n_{\_}M$ unique inputs in $\mathbf{X}_{\_}{n_{\_}M}$ when ignoring the $\xi$ column. In our case $n=n_{\_}M/31$. Let $\bar{\mathbf{X}}_{\_}n$ denote the matrix of those unique inputs, again without the $\xi$ column, so that $\bar{\mathbf{X}}_{\_}n$ is $n\times 4$. Now, consider each row $\bar{\mathbf{x}}_{\_}i$ of $\bar{\mathbf{X}}_{\_}n$, and let $y_{\_}j(\bar{\mathbf{x}}_{\_}i)$ denote the $j^{\mathrm{th}}$ replicate associated with $\bar{\mathbf{x}}_{\_}i$ among the $n_{\_}i$ replicates in $\mathbf{X}_{\_}{n_{\_}M}$. We have that all $n_{\_}i=31$. Then, calculate the first two moments for these replicates as
@@
Finally, store these in $\bar{\mathbf{Y}}_{\_}n$ and $\bar{\mathbf{S}}_{\_}n$, respectively, to line up with the rows of $\bar{\mathbf{X}}_{\_}n$. Note that we are dropping the $M$ subscripts with the understanding that we are only using this on the computer model runs. If the field data also had replicates we could additionally make use of them there, but our sensors record only one measurement per day.
SK involves modeling $\bar{\mathbf{Y}}_{\_}n$ and $\bar{\mathbf{S}}_{\_}n$ via GPs rather than operating directly on the original $D_{\_}{n_{\_}M}$. It turns out that these $2n$ quantities are not sufficient for $\mathbf{Y}_{\_}{n_{\_}M}$ but they are close. Binois, Gramacy and Ludkovski (2018) show that there are $12n+$ 1 sufficient statistics. Nevertheless, Ankenman, Nelson and Staum (2008) argue that inference based on these $2n$ quantities is unbiased asymptotically and minimizes mean-squared error. Now suppose we wish to predict $\bar{Y}$ with a GP. If the degree of replication is uniform, e.g., $n_{\_}i=31$ for all $i$, then we could simply plug the $\bar{\mathbf{Y}}_{\_}n$ values, along with $\bar{\mathbf{X}}_{\_}n$, into Eq. (4) after solving for MLE hyperparameters. We can do the same thing for the second moments $\bar{\mathbf{S}}_{\_}n$. This is the essence of SK with two caveats: (a) uniform replication is rare, although that is the case for NOAA-GLM; (b) it helps to explicitly link the mean and variance processes when making forecasts. First, fit a GP to $\bar{\mathbf{S}}_{\_}n$ on a transformed scale (Johnson et al., 2018) to respect positivity: $\mathcal{GP}(D^{(v)}_{\_}n)$ where $D^{(v)}_{\_}n=(\bar{\mathbf{X}}_{\_}n,T(\bar{\mathbf{S}}_{\_}n))$. For $T(\cdot)$ we prefer the square root to model standard deviations, although a logarithm also works well. Then feed $\mathcal{X}\equiv\bar{\mathbf{X}}_{\_}n$ into Eq. (4), and define
@@
Now consider new/forecasting coordinates $\mathcal{X}$ that may be of interest, like for a particular day $t$ and multiple horizons $h$ into the future, or representing hindcasts to compare to to sensor measurements discussed further in Section 3.4. Let $s(\mathcal{X})=\mathrm{Diag}(T^{-1}(\mu^{(v)}(\mathcal{X})))$ capture these back-transformed variance predictions – they are the predictive analog of Eq. (8). Using these, via $\mathbf{S}_{\_}n$ and $\mathcal{GP}(D^{(m)})$ where $D^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$, the SK equations are
$\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ $=\tau^{2}k(\bar{\mathbf{X}}_{\_}n,\mathcal{X})^{\top}(\tau^{2}k(\bar{\mathbf{X}}_{\_}n)+\mathbf{S}_{\_}n)^{-1}\boldsymbol{\bar{Y}}_{\_}{n}$ (9)
@@
Observe in Eq. (9) that augmenting the diagonal of $k(\bar{\mathbf{X}}_{\_}n)$ with $\mathbf{S}_{\_}n$ results in a scaled nugget-like term, by analogy to $\tau^{2}g$ in Eq. (4). It facilitates a signal-to-noise trade-off that is unique to each training data element. Smoothed (squared) standard errors $\mathbf{S}_{\_}n$ naturally downweight large variance/low replication training data. It may be shown that $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})=\mu_{\_}{n_{\_}M}(\mathcal{X})$ from Eq. (4), identically, which could represent a huge computational savings when $n\ll n_{\_}M$.
@@
Figure 5: 2d slices of predictive mean $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ (left) and variance $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}$ (middle) from the (heteroscedastic) GLM surrogate over depth (m) and horizon (days). Together, these summarize a hindcast for 2022-02-19. The right panel shows a 1m-depth slice with GLM ensemble forecasts (light gray lines), SK surrogate (darker, solid line) 90% CIs and PI (dotted and dashed lines). Observations are overlaid as darker colored dots.
@@
To illustrate, consider Figure 5 which is derived from fits of two GPs, for mean and variance, to a simulation campaign spanning three years over a thirty-day horizon. The data represented in the plot comprises a 2d slice of predictions from those two fits corresponding to a forecast beginning on 2022-02-19, offering a surrogate for the NOAA-GLM hindcast beginning on that date. Focusing first on the left panel, showing $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ from Eq. (9), observe that temperature is decreasing with increasing depth, and also generally increasing over horizon, most prominently at the surface (low depths). Both relationships are clearly nonlinear. Now looking at the middle panel providing $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}=\mathrm{Diag}(\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X}))$ we see a similar relationship in the spread of possible temperatures. Variance is much lower at lower depths, even at longer horizons, because these depths are more insulated from weather at the surface. The right panel of the figure shows a 1d slice of the surface(s) 1m depth along with additional information. Gray lines indicate raw NOAA-GLM simulations (i.e., a subset of $D_{\_}{n_{\_}M}$). Filled purple dots indicate sensor observations. These latter values are not part of the analysis at this time and are provided for reference only. Observe that the dashed orange 90% PIs, derived from $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ and $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}$, have appropriate coverage. Dotted lines indicate 90% CIs on the mean that are derived by eliminating the $s(\mathcal{X})$ term in Eq. (10), as an analogue to Eq. (5). It is this squared standard error which is crucial to our bias-correcting analysis described in Section 3.4. Finally, observe that NOAA-GLM has a cold bias compared to the field data.
@@ 3.3 GP surrogates and scalability
The surrogates behind the views in Figures 5 were constructed using the approximate methods we describe here. Scalable GP approximation has been an active field of late. Techniques include compactly supported kernels (Gneiting, 2002; Kaufman et al., 2011), fixed-rank kriging (Cressie and Johannesson, 2008) inducing points (Banerjee et al., 2008; Quinonero-Candela and Rasmussen, 2005; Hensman, Fusi and Lawrence, 2013) local divide-and-conquer (Emery, 2009; Gramacy, 2016) and the Vecchia/nearest-neighbor approximation (e.g., Datta et al., 2016; Katzfuss et al., 2020a) The so-called scaled Vecchia approximation (SVecchia; Katzfuss, Guinness and Lawrence, 2022) seems particularly well-suited to our setting because it allows coordinate-wise lengthscale estimation and scales in flops that are quasi-linear in $n$. Crucially, an open-source implementation is provided for R: https://github.com/katzfuss-group/scaledVecchia via GpGp (Guinness, 2018) and GPVecchia (Katzfuss et al., 2020b) packages. We provide a review of the Vecchia approximation, and some implementation details in Appendix A.1.
@@
Vecchia has never been used, and nor to our knowledge has any other large-scale GP approximation, to model input-dependent variances, like through SK (Section 3.2). Fitting second moments $\mathcal{GP}(D_{\_}n^{(v)}$) in order to get a predictor for $T(\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathcal{X}})\})$ is straightforward, because this amounts to feeding in (transformed) $s_{\_}i$-values (Eq. 7) as responses along with unique inputs $\bar{\mathbf{x}}_{\_}i$. The visual in the right panel of Figure 5 shows $\mu^{(v)}_{\_}n(\mathcal{X})$. Using Vecchia to fit the mean GP conditional on those variances (left panel of Figure 5) is harder because the R library we use does not accommodate a vector of nuggets, which comprises $\mathbf{S}_{\_}n$ in Eqs. (910). So we ended up implementing the following workaround instead.
@@
We begin by fitting first moments with an otherwise ordinary Vecchia approximated GP, i.e., $\mathcal{GP}(D_{\_}n^{(m)})$, where $D_{\_}n^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$. Let $\mu^{(m)}_{\_}n(\mathcal{X})$ denote the predictive mean of $\mathcal{GP}(D_{\_}n^{(m)})$. Note that $\mu^{(m)}_{\_}n\neq\mu^{\mathrm{SK}}_{\_}n$ from Eq. (9) because the former utilizes a covariance structure with a single nugget, i.e., $\tau^{2}(k(\bar{\mathbf{X}}_{\_}n)+g\mathbb{I}_{\_}n)$ following Eq. (4), and the latter uses $\tau^{2}k(\bar{\mathbf{X}}_{\_}n)+\mathbf{S}_{\_}n$. However, they are equal in expectation because GPs furnish unbiased predictors (see, e.g., Santner et al., 2018). Nevertheless, the two can be quite different if the degree of replication is high because of the $\mathbf{n}$-vector in the denominator of Eq. (8). Specifically, if the replication degree varies significantly from one location to another, for instance $n_{\_}i\gg n_{\_}j$ for $\bar{\mathbf{x}}_{\_}i$ and $\bar{\mathbf{x}}_{\_}j$, then the predictions from $\mu^{\mathrm{SK}}_{\_}n(\mathbf{x})$ for $\mathbf{x}$ in the vicinity of $\bar{\mathbf{x}}_{\_}j$, the site with fewer replicates, could be considerably more precise than $\mu^{(m)}_{\_}n(\mathbf{x})$, which does not know to place less trust in $\bar{y}_{\_}j$ than in other locations, such as near $\bar{\mathbf{x}}_{\_}i$. However, in our setting, we have uniform replication, with all $n_{\_}i=31$. In a study described in Appendix A.3, and summarized in Figure 13, we find there is essentially no difference between $\mu^{(m)}_{\_}n$, via Vecchia, and $\mu^{\mathrm{SK}}_{\_}n$.
@@
Approximating what SK provides for the variance of the first moment, $\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})$ in Eq. (10), with Vecchia, follows a similar thread. Whereas our development of the mean leverages the unbiased nature of Gaussian expectations, second moments follow standard calculations of Gaussian variances. Crucially, we do not need the entire covariance structure $\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})$ for bias correction in Section 3.4; we need only the diagonal. Those “variances of the mean” are Gaussian standard errors, i.e., $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}/n$, again leveraging uniform replication: $n_{\_}i=31$. Our Vecchia fit to second moments provides an estimate of $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}$ so all we need to do is divide this quantity by $n_{\_}i$, for any $i$, to get an estimate of the variance of the mean. However, we prefer a more conservative application of this in an SK-approximating context to avoid under-covering out-of-sample. We found that using the 95th quantile of the predictive mean obtained from $\mathcal{GP}(D_{\_}n^{(v)})$, notated as $\mu^{(v)95}_{\_}n(\mathbf{x})$, added to the variance of the first moments’ fit ($\sigma^{2m}_{\_}n(\mathbf{x})$) provided sufficient coverage. That is, we use
$\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathbf{x})$ $=\mu_{\_}n^{(m)}(\mathbf{x})$ and $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ $=\mu^{(v)95}_{\_}n(\mathbf{x})/n_{\_}i+\sigma^{2(m)}_{\_}n(\mathbf{x}).$ (11)